OpenStudy (anonymous):

Which graph is the solution to the inequality -3y-y+11>99

OpenStudy (anonymous):

@Akeller27 I think you can solve this one for yourself, do what you did in the other question.

OpenStudy (anonymous):

what about the -y i dont get that part

OpenStudy (saifoo.khan):

HINT: Add like terms.

OpenStudy (anonymous):

Well what is -3y-y?

OpenStudy (anonymous):

-4y > 110 y > -27.5

OpenStudy (anonymous):

am i right or wrong?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

Like I was saying before, when we have an inequality we solve just as we would if it was an equality. If the question was -3y-y+11=99 what would you do first?

OpenStudy (saifoo.khan):

@ramit ; the inequality sign switches when the variable gets rid of a negative sign.

OpenStudy (anonymous):

thanks for the info

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

subtract 11 from both sides so 99-11=88 and 11-11=0

OpenStudy (saifoo.khan):

@85295james : Direct answering on the site isn't allowed. :)

OpenStudy (anonymous):

Very good, but you also need to combine similar terms (hint: get the y terms together).

OpenStudy (anonymous):

then divide 99/-3=-29

OpenStudy (anonymous):

Have you gotten to the stage where you have -3y-y>88 Akeller?

OpenStudy (anonymous):

nevermind i figureed it out thanks anyways

OpenStudy (anonymous):

Even better :D

OpenStudy (anonymous):

thanks :)